Diego Company manufactures one product that is sold for $75 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 57,000 units and sold 52,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 25
Direct labor $ 18
Variable manufacturing overhead $ 3
Variable selling and administrative $ 5
Fixed costs per year:
Fixed manufacturing overhead $ 627,000
Fixed selling and administrative expense $ 645,000

The company sold 36,000 units in the East region and 16,000 units in the West region. It determined that $310,000 of its fixed selling and administrative expense is traceable to the West region, $260,000 is traceable to the East region, and the remaining $75,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

15. Assume the West region invests $47,000 in a new advertising campaign in Year 2 that increases its unit sales by 20%. If all else remains constant, what would be the profit impact of pursuing the advertising campaign?
Answer #1
Direct materials 25
Direct labor 18
Variable manufacturing overhead 3
Variable selling and administrative 5
Variable cost per unit 51
Contribution margin per unit = Selling price per unit - Variable cost per unit = 75 - 51 24
Increase in unit sales of west region = 16000 * 20% 3200
Increase (decrease) in profit = ( Increase in unit sales of west region * Contribution margin per unit ) - Cost of new advertising campaign = ( 3200 * 24 ) - 47000 29800
Profit will increase by $29,800